OpenStudy (anonymous):

At noon, ship A is 10 nautical miles due west of ship B. Ship A is sailing west at 18 knots and ship B is sailing north at 18 knots. How fast (in knots) is the distance between the ships changing at 7 PM? (Note: 1 knot is a speed of 1 nautical mile per hour.)

OpenStudy (radar):

I may be wrong, and if so, I hope someone will post to correct me. But it appears to me that ship A and B have a speed or rate of 18 knots, but as a velocity they are 90 degrees apart, so the distance change would be the square root of their sum:\[d` = \sqrt{18 + 18}\]

OpenStudy (radar):

That is all the question is asking for...

OpenStudy (phi):

radar has a nice insight, but the answer is ever so slightly different. let the distance between the two boats be D, and let x be the distance from the "origin" to boat A, and similarly , y the distance from the origin to boat B.  \[ D^2= x^2 + y^2 \] take the (implicit) derivative with respect to time t: \[ 2 D \dot{D} = 2 x \dot{x} + 2y \dot{y}\] \( \dot{x} \) is the speed of boat A = 18 nm/hr. and is equal to the speed of boat B. Thus \[ D \dot{D}= (x+y) 18 \\ \dot{D} = \frac{ 18(x+y) }{D} \] at 7 pm x is 10+7*18= 136 and y = 7*18= 126 use those values to find the velocity of D. We will get 25.437 which is very close to radar's \[ \sqrt{18^2 +18^2} = 25.456 \]

OpenStudy (radar):

Thanks @phi I had my doubts, somehow I figured those distances had to be brought in, and I did try to do the related rate derivative but I went astray and got an answer which was over twice their individual rate.....so I knew that was wrong, so I took the easy way out. Thanks for refreshing my memory on the proper steps.